i dont understand how to solve this: e^2x-e^x-30=0
Answer by edjones(8007)

e^2x-e^x-30=0
(e^x-6)(e^x+5)=0
e^x=6
ln(e^x)=ln(6)
x=ln^(6)
.
Ed